OpenStudy (anonymous):

How do I simplify sinx(1/cosx) ?

OpenStudy (anonymous):

lol sin x/cos x = tan x :D

OpenStudy (anonymous):

how did you get that? :o

OpenStudy (anonymous):

sin x = altitude/hypotenuse cos x = base/hypotenuse sinx/cosx = (altitude/hypotenuse)/(base/hypotenuse) = altitude/base (hypotenuse cancel each other) which is out own tan x :P

OpenStudy (anonymous):

our*

OpenStudy (anonymous):

so will (cosx)(tanx) be (cosx/tanx)(sinx/cosx) ?

OpenStudy (anonymous):

cos x tan x = cos x (sin x /cos x) = sin x

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

what about 1-cos^2x ? will it be sin^2x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and 1-sin^2x will be cos^2x? wow thank you :)

OpenStudy (anonymous):

sin^2x+cos^2x = 1 sec^2x = 1 + tan^2x cosec^2x = 1 + cot^2x (Most important formulae) :P

OpenStudy (anonymous):

oh okay yes. um what wiill cos^2x+sin^2x be? would I do this, 1-sin^2x+1-cos^2x ? which would be -sin^2x-cos^2x ?

OpenStudy (anonymous):

first of all.. x + y = y + x -sin^2x-cos^2x = -(sin^2x+cos^2x) = -1

OpenStudy (anonymous):

sin^2x+cos^2x = 1 = cos^2x+sin^2x

OpenStudy (anonymous):

ohh yes. omg thank you sooo much ! I am very grateful :)

OpenStudy (anonymous):

Happy to help :))

OpenStudy (anonymous):

(1-sinx)(1+sinx) would be 1^2-sinx^2 ?

OpenStudy (anonymous):

do I expand or replace the sinx with tanx/cosx?

OpenStudy (anonymous):

why would you replace.. first one is correct

OpenStudy (anonymous):

which is also cos^2x

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

for tanx/sinx would I do sinx/cosx times tanx/cosx?

OpenStudy (anonymous):

then sinxtanx/cosx^2 ? :S

OpenStudy (anonymous):

why are you making it so complicated.. be cool :P

OpenStudy (anonymous):

then what should I do? :(

OpenStudy (anonymous):

remember tan x = sinx/cosx cot x = cosx/sinx

OpenStudy (anonymous):

so if it is tanx/sinx dont I put sinx/cox for tanx ? replace it? and for sinx I would puttanx/cosx? :S Im confused

OpenStudy (anonymous):

yes u can do that.. remember that if it is an identity u can replace everything to simplify

OpenStudy (anonymous):

in tanx/sinx .. if u put tanx=sinx/cosx... sinx will cancel each other and u get 1/cosx

OpenStudy (anonymous):

which is also sec x =)

OpenStudy (anonymous):

hmm okayy. well for 1+tan^2x/tan^2x, will I replace the tan^2x with sec^2x ?

OpenStudy (anonymous):

noooo.. replace 1+tan^2x with sec^2x NOT the tan^2x alone

OpenStudy (anonymous):

OH okay, then what ?

OpenStudy (anonymous):

what "then what"

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

well after i sub in, 1+sec^2x-1/tan^2x what do I do ? the 1 cancels out right? then i have sec^2x/tan^2x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay thannn what should I do? becaue im suppose to get 1/sin^2x :(

OpenStudy (anonymous):

sec^2x/tan^2x = cos^2x/sin^2x.cos^2x cos^2x cancel.. and 1/sin^2x remains

OpenStudy (anonymous):

achaaa, okay thank you soooo much !

OpenStudy (anonymous):

for sinxcosx/1-sin^2x, i would do sinxcox/1, then would I replace sinx or cosx?

OpenStudy (anonymous):

the answer is sinx cosx .. nothing to simplify here

OpenStudy (anonymous):

but in my book it says the answer is tanx

OpenStudy (anonymous):

oops.. lol.. my mistake its sinxcosx/cos^2x = sinx/cosx = "the answer"

OpenStudy (anonymous):

what happenes to the cos^2x?

OpenStudy (anonymous):

X/X^2 = ??

OpenStudy (anonymous):

x ?

OpenStudy (anonymous):

oh nvm I get it ! :D

OpenStudy (anonymous):

so 1-cos^2x it says factor. it says (1+cosx)(1-cosx) but why can't it be sin^2x?

OpenStudy (anonymous):

it can be..

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

convert in a way so that it can be simplified..

OpenStudy (anonymous):

alright, thanks man !

OpenStudy (anonymous):

it comes with practice.. (solved examples also)

OpenStudy (anonymous):

aha yeah it does. I need alot since I just started